The mean weight of 500 college students is 70 kg and the standard deviation is 5 kg

Assuming that the weight is normally distributed, determine how many students weights:

a: Between 60 kg and 75 kg

b. More Than 90 kg

c. less than 65 kg



1
Expert's answer
2021-03-29T01:44:21-0400

By condition, "a = 70\,\,\sigma = 5"

a: Between 60 kg and 75 kg

Find the probability that the student's weight will be between 60 and 75 kg:

"P(60 < x < 75) = \Phi \left( {\frac{{\beta - a}}{\sigma }} \right) - \Phi \left( {\frac{{\alpha - a}}{\sigma }} \right) = \Phi \left( {\frac{{75 - 70}}{5}} \right) - \Phi \left( {\frac{{60 - 70}}{5}} \right) = \Phi \left( 1 \right) + \Phi \left( 2 \right) = 0,3413 + 0,4772 = {\rm{0}}{\rm{,8185}}"

Then quantity of students weights between 60 kg and 75 kg is

"{\rm{500}} \cdot {\rm{0}}{\rm{,8185}} \approx {\rm{409}}"

Answer: 409

b. More Than 90 kg

"P(x > 90) = \Phi \left( \infty \right) - \Phi \left( {\frac{{90 - 70}}{5}} \right) = \Phi \left( \infty \right) - \Phi \left( 4 \right)\approx 0,5 - 0,5 \approx 0"

Then quantity of students weights more then 90 kg is 0.

Answer: 0

c. less than 65 kg

"P(x < 65) = P(0 < x < 65) = \Phi \left( {\frac{{65 - 70}}{5}} \right) - \Phi \left( {\frac{{0 - 70}}{5}} \right) = \Phi \left( {14} \right) - \Phi \left( 1 \right) = 0,5 - 0,3413= {\rm{0}}{\rm{,1587}}"

Then quantity of students weights less then 65 kg is "{\rm{500}} \cdot {\rm{0}}{\rm{,1587}} \approx 79"

Answer: 79